What is the length of side A, in the triangle of A = 25°, b = 28, a = 18?
Answer by Herdaysoji(25)

First sine rule; (sinA/a)=(sinB/b) cross multiplication
sinB=(sinA*b)/a=(sin25*28)/18=(0.4226..*28)/18
sinB=11.8333/18=0.6574
B=(sin^-1)0.6574=41.1
.'.A+B+C=180.'.28+41+C=180
C=180-28-41=111
using cosine rule to get c
(c^2)=(a^2)+(b^2)-2ab(cosC)
(c^2)=(18^2)+(28^2)-2*18*28 (cos111)
(c^2)=324+784-(1008*-0.3584)
(c^2)=1108-(-3.61.234.....)
(c^2)=1108+361.234....
(c^2)=1469.234..
c=square root(1469.234....)
c=38.330.....